# An Uncertainty Quantification Framework for Deep Learning-Based Automatic Modulation Classification

### v1

@@ Abstract
Deep learning has been shown to be highly effective for automatic modification classification (AMC), which is a pivotal technology for next-generation cognitive communications. Yet, existing deep learning methods for AMC often lack robust mechanisms for uncertainty quantification (UQ). This limitation restricts their ability to produce accurate and reliable predictions in real-world environments, where signals can be perturbed as a result of several factors such as interference and low signal-to-noise ratios (SNR). To address this problem, we propose a deep ensemble approach that leverages multiple convolutional neural networks (CNNs) to generate predictive distributions, as opposed to point estimates produced by standard deep learning models, which produce statistical characteristics that quantify the uncertainty associated with each prediction. We validate our approach using real-world AMC data, evaluating performance through multiple UQ metrics in a variety of signal environments. Our results show that our proposed ensemble-based framework captures uncertainty to a greater degree compared to previously proposed baselines in multiple settings, including in-distribution samples, out-of-distribution samples, and low SNR signals. These findings highlight the strong UQ capabilities of our ensemble-based AMC approach, paving the way for more robust deep learning-based AMC.
Index Terms:
Automatic modulation classification, deep ensembles, deep learning, out-of-distribution samples, uncertainty quantification,
@@ I Introduction
Automatic Modulation Classification (AMC) is a vital component of modern wireless networks, as it enables the identification of modulation schemes without prior knowledge of the transmitted signal. This capability is crucial for ensuring reliable communication in dynamic and noisy environments, including military [1, 2] and civilian applications [3], where signals can be subject to interference and environmental distortions. As the Internet of Things (IoT) expands, next-generation communications (e.g., 6G) will heavily rely on AMC for efficient management of the increasingly congested wireless spectrum [4, 5].
@@
Traditional AMC methods, such as maximum likelihood-based (MLB) approaches [6, 7, 8, 9], rely on statistical tests to estimate the probability of a particular modulation scheme. Although these methods can be effective, they require substantial prior knowledge about the received signal and its channel conditions [10]. Furthermore, MLB methods tend to be computationally expensive, incurring high latencies while requiring manual feature engineering from received in-phase and quadrature (IQ) time samples. Such limitations restrict their scalability and adaptability into the high-volume IoT spectrum in real-world scenarios.
@@
To address these challenges deep learning (DL) has recently emerged as a powerful alternative to MLB methods for AMC. DL models can efficiently learn distinguishing AMC characteristics from received IQ signals without requiring manual feature engineering while achieving state-of-the-art classification accuracy [11, 3, 12, 13, 14]. As a result, they provide a promising alternative to MLB approaches to meet the demands of overcrowded next-generation cellular networks. However, despite their success, DL models exhibit a critical limitation. Specifically, deep learning-based AMC systems are frequently overconfident [15] in their predictions, even when they are incorrect, and often lack robust mechanisms for uncertainty quantification (UQ) to characterize the confidence associated with a given prediction. This overconfidence can result in significant performance degradations during deployment, particularly on incorrect predictions, when the model encounters previously unseen signal environments. Such overconfidence is further exacerbated in mission-critical settings, where decisions must be made under uncertainty, including low SNR conditions and adversarial scenarios [16, 17].
@@
To address these challenges, we propose a UQ framework for AMC based on an ensemble of convolutional neural networks (CNNs), which generate predictive distributions, as opposed to point estimates produced by current state-of-the-art CNNs, whose statistical characteristics quantify the uncertainty associated with each prediction. Building on ensemble learning techniques [18, 19, 20, 21, 22], our approach provides calibrated uncertainty estimates, thereby enhancing reliability and robustness of current DL-based AMC frameworks. Compared to traditional UQ methods such as Bayesian Neural Networks (BNNs) [23], our proposed framework offers superior performance with lower computational overhead. Specifically, our method achieves higher classification accuracy in multiple signal environments, demonstrates better performance on multiple UQ metrics, and displays enhanced resilience on out of distribution (OOD) samples such as adversarial examples [24, 25, 26]. Moreover, our ensemble-based approach is scalable to large-parameter architectures such as ResNet [27], enabling existing DL-based AMC frameworks to improve their performance while simultaneously incorporating UQ for increased robustness.
@@ II Related Works
Earlier methods for AMC relied on likelihood methods [10, 7, 9], which provided theoretical guarantees as well as uncertainty estimation. However, such methods incur computational inefficiency in large-scale networks, making them difficult to adopt in the increasingly congested wireless spectrum. DL, as an alternative to MLB approaches, has gained significant attention for AMC due to its state-of-the-art classification performance and significantly lower inference latencies compared to MLB approaches. In this regard, various neural network architectures have been explored such as CNNs [28, 3, 4, 29], recurrent neural networks (RNNs) [30, 31, 32], and long short-term memory (LSTM) [33]. By learning hierarchical features directly from raw signals, these DL approaches, in addition to having low online latencies, often excel in AMC classification tasks with little to no need for feature engineering. However, while prior studies have demonstrated the efficiency of DL compared to MLB approaches for AMC, less attention has been given to the UQ capabilities of these methods, which is the focus of our work.
More recently, Bayesian neural networks (BNNs) [34] have been proposed to perform DL-based AMC with UQ capabilities. Although BNNs seemingly provide the advantage of both the high classification performance of DL as well as UQ capabilities, their practical application is hindered by lower classification performance compared to state-of-the-art CNNs. Despite their ability to capture predictive distribution and variance, which aids in uncertainty estimation, BNNs often have significantly lower classification performance compared to CNNs for AMC. This gap in performance and adaptability motivates our development of ensemble-based UQ methods for AMC.
Ensemble learning has emerged as a compelling alternative to BNNs in multiple domains, such as image processing [19] and bioinformatics [35], where multiple models are jointly analyzed. Compared to single models, deep ensembles yield superior classification performance [36] and generate predictive distributions instead of point estimates, leading to well-calibrated UQ estimates [19, 37, 21]. Additionally, deep ensembles have demonstrated higher classification accuracy in the presence of data distortions such as OOD samples, which have been shown to limit the efficacy of DL-based AMC methods [24, 25, 26].
@@ III Methodology
In this section, we develop our deep ensemble-based AMC framework and describe the metrics used to quantify the uncertainty of each model. We begin by discussing our AMC signal model in Sec. III-A. We then present our deep ensemble approach in Sec. III-B. Subsequently, Sec. III-C details the estimation metrics employed to evaluate the performance of different methods. Finally, Sec. III-D discusses our approach for analyzing OOD AMC signals.
@@ III-A Signal Modeling
where $\mathbf{H}=\text{diag}\{h[0],\cdots,h[\ell-1]\}\in\mathbb{C}^{\ell\times\ell}$, $\mathbf{n}\in\mathbb{C}^{\ell}$ represents additive white Gaussian noise (AWGN), and $\rho$ denotes the signal to noise ratio (SNR), which is typically unknown at receiver further motivating the need for a UQ-based AMC framework. Moreover, we map each received baseband signal to a two-dimensional real matrix, $\mathbf{r}\in\mathbb{C}^{\ell}\rightarrow\mathbf{r}\in\mathbb{R}^{\ell\times 2}$ where the first and second column of $\mathbf{r}$ represent the in-phase and quadrature components, respectively, of $\mathbf{r}$ for compatibility with real-valued neural networks.
The receiver aims to perform AMC by calculating $\text{argmax}_{i}P(m_{i}|\mathbf{r},\theta)$, where $\theta$ parameterizes the model used to calculate the AMC probability (further discussed in Sec. III-B), $m_{i}\in\mathcal{M}$, and =$\mathcal{M}=\{m_{1},\cdots,m_{C}\}$ represents the set of considered modulation constellations. We assume that the receiver uses $\mathcal{X}_{\text{tr}}=\{\mathbf{r}(n),\mathbf{y}(n);n=1,\cdots,N\}$ and $\mathcal{X}_{\text{te}}=\{\mathbf{r}(t),\mathbf{y}(t);t=1,\cdots,T\}$ as the training and testing datasets, respectively, where $\mathcal{X}_{\text{tr}}\cap\mathcal{X}_{\text{te}}=\emptyset$.
@@ III-B Ensemble Modeling
Data-driven AMC models, alone, cannot effectively characterize the uncertainty associated with their predictions because they tend to produce overconfident outputs. To address this limitation, we adopt an ensemble modeling approach consisting of multiple state-of-the-art deep learning AMC classifiers [38]. Unlike a standalone model, ensemble predictions cannot be obtained directly because multiple models are used simultaneously. To ensure diversity in predictions, each classifier is initialized with random parameters while maintaining the same architecture, resulting in each model converging to different local minima thus ensuring parallelized training as well as diverse predictions despite training on the same dataset [18]. As a result, the ensemble exhibits variability, yielding varying levels of confidence on samples that are known to be difficult for deep learning classifiers to operate on such as low SNR signals. As a result, analyzing the joint output prediction from each classifier in the ensemble simultaneously improves the UQ of the model.
@@
where $\text{argmax}_{i}P(m_{i}|\mathbf{r},\theta)$ yields the predicted modulation constellation from the ensemble. In addition to obtaining the probability with which the input belongs to a particular modulation classification, we use the distributive estimate produced by the ensemble to characterize additional UQ metrics, which are discussed in Sec. III-C.
III-C Estimation Metrics
Beyond obtaining the modulation classification estimate given in (4), we also employ several proper UQ scoring metrics [19] to assess model performance comprehensively. In this capacity, we consider the negative log-likelihood (NLL), the Brier score, the width of the prediction’s confidence interval (CI), the coverage of the prediction, and set of high-confidence predictions.
@@ III-C Estimation Metrics
$-\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{C}\mathbf{y}_{tj}\log(\hat{\mathbf{y}}_{tj}),$ (5)
where lower NLL values indicate that the model assigns higher confidence to the correct classes, reflecting better alignment between predictions and true labels. Similarly, the Brier score, which measures the mean squared difference between the predicted probabilities and the actual outcomes, is computed over all samples and classes and is given by
@@
where lower Brier score values indicate that the model’s probabilistic predictions are closer to ground truth values, minimizing deviations.
@@
$z_{\alpha}$ is the $21-\alpha/$ 2 quantile of a zero-mean unit variance Gaussian distribution, and $S^{2}$ is the variance of the predicted probability among models, and the CI width is given by the computing the difference between the upper and lower bounds of (7).
@@
We next calculate the coverage proportion, which evaluates the number of samples in the testing set whose true label is within the $z_{\alpha}$ CI. In computing the coverage of the ensemble, we consider two scenarios. First, we employ a strict condition, where the CI of the true class must contain one and the CI of all other classes must contain zero. This scenario strictly requires the ensemble to be certain that the predicted sample simultaneously (i) is modulated according to the predicted class and (ii) is not modulated according to any other class. Second, we relax these conditions and only require the CI of the true class to contain one and we do not consider the CIs of the other classes. This metric, in both scenarios, quantifies the proportion of samples that the ensemble is expected to be confident in. This scenario is particularly helpful for low SNR signals, which often suffer from lower classification performance in data-driven AMC approaches.
@@
Lastly, we define the high confidence set to evaluate the frequency with which the ensemble is highly confident in its predictions. A high confidence prediction occurs when $\text{argmax}_{i}P(m_{i}|\mathbf{r},\theta)>0.8$. A high confidence prediction indicates overconfidence, which often occurs in singular deep learning models, while a low confidence prediction reflects under-confidence. Ideally, the ensemble should strike a balance, avoiding both extremes.
III-D Out-of-distribution Samples

### v2

@@ Abstract
Deep learning has been shown to be highly effective for automatic modulation classification (AMC), which is a pivotal technology for next-generation cognitive communications. Yet, existing deep learning methods for AMC often lack robust mechanisms for uncertainty quantification (UQ). This limitation restricts their ability to produce accurate and reliable predictions in real-world environments, where signals can be perturbed as a result of several factors such as interference and low signal-to-noise ratios (SNR). To address this problem, we propose a deep ensemble approach that leverages multiple convolutional neural networks (CNNs) to generate predictive distributions, as opposed to point estimates produced by standard deep learning models, which produce statistical characteristics that quantify the uncertainty associated with each prediction. We validate our approach using real-world AMC data, evaluating performance through multiple UQ metrics in a variety of signal environments. Our results show that our proposed ensemble-based framework captures uncertainty to a greater degree compared to previously proposed baselines in multiple settings, including in-distribution samples, out-of-distribution samples, and low SNR signals. These findings highlight the strong UQ capabilities of our ensemble-based AMC approach, paving the way for more robust deep learning-based AMC.
@@ I Introduction
Automatic Modulation Classification (AMC) is a vital component of modern wireless networks, as it enables the identification of modulation schemes without prior knowledge of the transmitted signal. This capability is crucial for ensuring reliable communication in dynamic and noisy environments, including military [mil_apps1, mil_apps2] and civilian applications [ref6], where signals can be subject to interference and environmental distortions. As the Internet of Things (IoT) expands, next-generation communications (e.g., 6G) will heavily rely on AMC for efficient management of the increasingly congested wireless spectrum [ref50, amc_survey].
@@
Traditional AMC methods, such as maximum likelihood-based (MLB) approaches [ref3, ref33, ref34, ref35], rely on statistical tests to estimate the probability of a particular modulation scheme. Although these methods can be effective, they require substantial prior knowledge of the received signal and its channel conditions [ref29]. Furthermore, MLB methods tend to be computationally expensive, incurring high latencies while requiring manual feature engineering from received in-phase and quadrature (IQ) time samples. Such limitations restrict their scalability and adaptability into the high-volume IoT spectrum in real-world scenarios.
@@
To address these limitations, deep learning (DL) has recently emerged as a powerful alternative to MLB methods for AMC. DL models can efficiently learn distinguishing AMC characteristics from received IQ signals without requiring manual feature engineering while achieving state-of-the-art classification accuracy [ref32, ref6, ref24, ref25, ref49]. As a result, they provide a promising alternative to MLB approaches to meet the demands of overcrowded next-generation cellular networks. However, despite their success, DL models exhibit a critical limitation. Specifically, deep learning-based AMC systems are frequently overconfident [ref8] in their predictions, even when they are incorrect, and often lack robust mechanisms for uncertainty quantification (UQ) to characterize the confidence associated with a given prediction. This overconfidence can result in significant performance degradations during deployment, particularly on incorrect predictions, when the model encounters previously unseen signal environments. Such overconfidence is further exacerbated in mission-critical settings, where decisions must be made under uncertainty, including low SNR conditions and adversarial scenarios [ref53, ref55].
@@
To address these challenges, we propose a UQ framework for AMC based on an ensemble of convolutional neural networks (CNNs), which generate predictive distributions, as opposed to point estimates produced by current state-of-the-art CNNs, whose statistical characteristics quantify the uncertainty associated with each prediction. Building on ensemble learning techniques [ref11, ref12, ref15, ref18, ref19], our approach provides calibrated uncertainty estimates, thereby enhancing reliability and robustness of current DL-based AMC frameworks. Compared to traditional UQ methods such as Bayesian Neural Networks (BNNs) [ref36], our proposed framework offers superior performance with lower computational overhead. Specifically, our method achieves higher classification accuracy in multiple signal environments, demonstrates better performance on multiple UQ metrics, and displays enhanced resilience on out of distribution (OOD) samples such as adversarial examples [amc_adv_atk, amc_adv_atk2, ref40]. Moreover, our ensemble-based approach is scalable to large-parameter architectures such as ResNet [ref37], enabling existing DL-based AMC frameworks to improve their performance while simultaneously incorporating UQ for increased robustness.
To our knowledge, this is the first work to evaluate the efficacy of deep ensembles for AMC. While ensemble techniques have been used in deep learning, their application in AMC poses unique challenges due to domain-specific distortions (e.g., dynamic SNR [ref32] and adversarial AMC conditions [amc_adv_atk]), and our results demonstrate that equal-weighted ensembles outperform more complex methods like BNNs and SNR-aware weighted ensembles in both classification and uncertainty estimation.
@@ II Related Works
Earlier methods for AMC relied on likelihood methods [ref29, ref33, ref35], which provided theoretical guarantees as well as uncertainty estimation. However, such methods incur computational inefficiency in large-scale networks, making them difficult to adopt in the increasingly congested wireless spectrum. DL, as an alternative to MLB approaches, has gained significant attention for AMC due to its state-of-the-art classification performance and significantly lower inference latencies compared to MLB approaches. In this regard, various neural network architectures have been explored such as CNNs [ref5, ref6, ref50, ref9], recurrent neural networks (RNNs) [ref42, ref43, ref44], and long short-term memory (LSTM) [ref41]. By learning hierarchical features directly from raw signals, these DL approaches, in addition to having low online latencies, often excel in AMC classification tasks with little to no need for feature engineering. However, while prior studies have demonstrated the efficiency of DL compared to MLB approaches for AMC, less attention has been given to the UQ capabilities of these methods, which is the focus of our work.
More recently, Bayesian neural networks (BNNs) [ref2] have been proposed to perform DL-based AMC with UQ capabilities. Although BNNs seemingly provide the advantage of both the high classification performance of DL as well as UQ capabilities, their practical application is hindered by lower classification performance compared to state-of-the-art CNNs. Despite their ability to capture predictive distribution and variance, which aids in uncertainty estimation, BNNs often have significantly lower classification performance compared to CNNs for AMC. This gap in performance and adaptability motivates our development of ensemble-based UQ methods for AMC.
Ensemble learning has emerged as a compelling alternative to BNNs in multiple domains, such as image processing [ref12] and bioinformatics [de_bi], where multiple models are jointly analyzed. Compared to single models, deep ensembles yield superior classification performance [ref45] and generate predictive distributions instead of point estimates, leading to well-calibrated UQ estimates [ref12, ref46, ref18]. Additionally, deep ensembles have demonstrated higher classification accuracy in the presence of data distortions such as OOD samples, which have been shown to limit the efficacy of DL-based AMC methods [amc_adv_atk, amc_adv_atk2, ref40].
@@ III Methodology
In this section, we develop our deep ensemble-based AMC framework and describe the metrics used to quantify the uncertainty of each model. We begin by discussing our AMC signal model in Sec. III-A. We then present our deep ensemble approach in Sec. III-B and its complexity analysis in Sec. III-C. Subsequently, Sec. III-D details the estimation metrics employed to evaluate the performance of different methods. Finally, Sec. III-E discusses our approach for analyzing OOD AMC signals.
@@ III-A Signal Modeling
where $\mathbf{H}=\text{diag}\{h[0],\cdots,h[\ell-1]\}\in\mathbb{C}^{\ell\times\ell}$, $\mathbf{n}\in\mathbb{C}^{\ell}$ represents additive white Gaussian noise (AWGN), and $\rho$ denotes the signal to noise ratio (SNR) of the received signal. Furthermore, we map each received baseband signal to a two-dimensional real matrix, $\mathbf{r}\in\mathbb{C}^{\ell}\rightarrow\mathbf{r}\in\mathbb{R}^{\ell\times 2}$ where the first and second column of $\mathbf{r}$ represent the in-phase and quadrature components, respectively, of $\mathbf{r}$ for compatibility with real-valued neural networks.
The receiver aims to perform AMC by calculating $\text{argmax}_{i}P(m_{i}|\mathbf{r},\theta)$, where $\theta$ parameterizes the model used to calculate the AMC probability (further discussed in Sec. III-B), $m_{i}\in\mathcal{M}$, and $\mathcal{M}=\{m_{1},\cdots,m_{C}\}$ represents the set of considered modulation constellations. We assume that the receiver uses $\mathcal{X}_{\text{tr}}=\{\mathbf{r}(n),\mathbf{y}(n);n=1,\cdots,N\}$ and $\mathcal{X}_{\text{te}}=\{\mathbf{r}(t),\mathbf{y}(t);t=1,\cdots,T\}$ as the training and testing datasets, respectively, where $\mathcal{X}_{\text{tr}}\cap\mathcal{X}_{\text{te}}=\emptyset$.
@@ III-B Ensemble Modeling
Data-driven AMC models, alone, cannot effectively characterize the uncertainty associated with their predictions because they tend to produce overconfident outputs. To address this limitation, we adopt an ensemble modeling approach consisting of multiple state-of-the-art deep learning AMC classifiers [ref52]. Unlike a standalone model, ensemble predictions cannot be obtained directly because multiple models are used simultaneously. To ensure diversity in predictions, each classifier is initialized with random parameters while maintaining the same architecture, resulting in each model converging to different local minima thus ensuring parallelized training as well as diverse predictions despite training on the same dataset [ref11]. As a result, the ensemble exhibits variability, yielding varying levels of confidence on samples that are known to be difficult for deep learning classifiers to operate on such as low SNR signals. As a result, analyzing the joint output prediction from each classifier in the ensemble simultaneously improves the UQ of the model.
@@
where $\text{argmax}_{i}P(m_{i}|\mathbf{r},\theta)$ yields the predicted modulation constellation from the ensemble. In addition to obtaining the probability with which the input belongs to a particular modulation classification, we use the distributive estimate produced by the ensemble to characterize additional UQ metrics, which are discussed in Sec. III-D.
III-C Complexity Analysis
Here, we analyze the computational complexity of our proposed framework during inference. The computational complexity of a single layer, $l$, of a CNN during runtime is given by $\mathcal{O}(C_{\text{in}}^{(l)}\cdot C_{\text{out}}^{(l)}\cdot K_{w}^{(l)}\cdot K_{h}^{(l)}\cdot H^{(l)}\cdot W^{(l)})$ and, thus, the complexity over an entire CNN is given by the sum of these complexities, resulting in
$\sum_{l=1}^{L}\mathcal{O}\bigg(C_{\text{in}}^{(l)}\cdot C_{\text{out}}^{(l)}\cdot K_{w}^{(l)}\cdot K_{h}^{(l)}\cdot H^{(l)}\cdot W^{(l)}\bigg),$ (5)
where $L$ is the number of layers and, in each layer $l$, $C_{\text{in}}^{(l)}$ is the number of input channels, $C_{\text{out}}^{(l)}$ is the number of output channels, $K_{w}^{(l)}$ and $K_{h}^{(l)}$ are the width and height of each convolutional kernel, and $H^{(l)}$ and $W^{(l)}$ represent the output spatial size. Since each CNN in our proposed framework is independent, they can process each received signal in parallel and thus operate without incurring any additional computational overhead compared to a single CNN. Thus, the computational complexity of our proposed method is given by (5). Moreover, in terms of wall clock time, each sample requires, on average, 1 ms per sample to process. Thus, our proposed framework does not incur any additional overhead in comparison to state-of-the-art deep learning-based AMC approaches.
III-D Estimation Metrics
Beyond obtaining the modulation classification estimate given in (4), we also employ several proper UQ scoring metrics [ref12] to assess model performance comprehensively. In this capacity, we consider the negative log-likelihood (NLL), the Brier score, the width of the prediction’s confidence interval (CI), the coverage of the prediction, the set of high-confidence predictions, the expected calibration error (ECE), and the kullback-leibler (KL) divergence.
@@ III-D Estimation Metrics
$\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{C}\mathbf{y}_{tj}\log(\hat{\mathbf{y}}_{tj}),$ (6)
where lower NLL values indicate that the model assigns higher confidence to the correct classes. In AMC, low NLL is important because poor model training under low SNR conditions can lead to overconfident misclassifications. Thus, NLL serves as an indicator of probabilistic robustness under channel impairments.
Similarly, the Brier score, which measures the mean squared difference between the predicted probabilities and the actual outcomes, is computed over all samples and classes and is given by
@@
where lower Brier score values indicate that the model’s probabilistic predictions are closer to ground truth values, minimizing deviations. Unlike accuracy, the Brier score penalizes both overconfidence in incorrect predictions and under confidence in correct predictions. In AMC systems, where prediction certainty impacts communication reliability, the Brier score provides an interpretable measure of prediction quality.
@@
$z_{\alpha}$ is the $1-\alpha/2$ quantile of a zero-mean unit variance Gaussian distribution, and $S^{2}$ is the variance of the predicted probability among models, and the CI width is given by computing the difference between the upper and lower bounds of (8). In AMC applications, CI widths offer a direct measure of the ensemble’s epistemic uncertainty, resulting in a quantified interpretable measure of the confidence associated with each signal’s modulation constellation prediction.
@@
We next calculate the coverage proportion, which evaluates the number of samples in the testing set whose true label is within the $z_{\alpha}$ CI. In computing the coverage of the ensemble, we consider two scenarios. First, we employ a strict condition, where the CI of the true class must contain one and the CI of all other classes must contain zero. This scenario strictly requires the ensemble to be certain that the predicted sample simultaneously (i) is modulated according to the predicted class and (ii) is not modulated according to any other class. Second, we relax the second condition and only require the CI of the true class to contain one and we do not consider the CIs of the other classes. This metric, in both scenarios, quantifies the proportion of samples for which the ensemble is confident at a $z_{\alpha}$ level that the input belongs to the predicted class. The coverage is especially helpful for understanding the ensemble’s confidence under low SNR conditions, which often is the cause of lowered classification performance in AMC applications. It is important to note that such confidence intervals can extend beyond the $[0,1]$ range particularly when the mean is near zero or one, indicating that the model is generally confident in its prediction about the correct class while simultaneously being confident that the sample does not belong to another class (i.e., the strict condition) or merely being confident in its prediction about the correct class without considering the confidence of other incorrect classes (i.e, the relaxed condition).
@@
To explore confidence further, we define the high confidence set to evaluate the frequency with which the ensemble is highly confident in its predictions. A high confidence prediction occurs when $\text{argmax}_{i}P(m_{i}|\mathbf{r},\theta)>0.8$. A high confidence prediction indicates overconfidence, which often occurs in singular deep learning models, while a low confidence prediction reflects under-confidence. Ideally, in AMC applications, the ensemble should strike a balance, avoiding both extremes. This metric, like others previously discussed, provides a measure for model reliability during adverse channel conditions in AMC applications.
We also compute the Expected Calibration Error (ECE), which measures how far the model’s predicted probabilities are from the true frequency of correct predictions. The ECE is given by
$\sum_{k=1}^{\mathcal{K}}\frac{\lvert\zeta_{k}\rvert}{T}\;\bigl|\mathrm{acc}(\zeta_{k})-\mathrm{conf}(\zeta_{k})\bigr|,$ (9)
where, for each bin $\zeta_{k}$ we take the fraction of samples in that bin, $\lvert\zeta_{k}\rvert/T$, multiply it by the absolute difference between the bin’s actual accuracy $\mathrm{acc}(\zeta_{k})$ and its average predicted confidence $\mathrm{conf}(\zeta_{k})$, and then sum these values over all $\mathcal{K}$ bins to yield the ECE. A lower ECE indicates that the model’s prediction estimates are well‐calibrated to its actual accuracy.
Finally, we measure the Kullback–Leibler (KL) divergence between the one-hot true distribution $\mathbf{y}_{t}$ and the predictive distribution $\hat{\mathbf{y}}_{t}$ for each sample, the KL divergence is given by
$D_{\mathrm{KL}}\bigl(\mathbf{y}_{t}\,\big\|\,\hat{\mathbf{y}}_{t}\bigr)\;=\;\sum_{j=1}^{C}y_{tj}\,\log\!\Bigl(\tfrac{y_{tj}}{\hat{y}_{tj}}\Bigr).$ (10)
Here, lower KL divergence indicates that the model assigns higher probability to the true modulation class, yielding more accurate distributions. In AMC, this reflects both the confidence and correctness of probabilistic predictions under noisy channel conditions.
III-E Out-of-distribution Samples